Leo is 38 years older than Adam. Five times Adam`s age, when decreased by 5 is 1 more than Leo`s age. How old is each.
Answer by ichudov(499)

Rewrite the problem to look more mathematical.
Leo = Adam + 38
5*Adam-5 = Leo+1
Because Leo=Adam+38,
5*Adam-5=Adam+38+1
4*Adam=38+1+5
4*Adam = 44
Answer:
Adam = 11
Leo = 11+38 = 49